eBook Coastal Carolina Heating and Cooling Inc. has a 6-month backlog of orders for its patented...

eBook

Coastal Carolina Heating and Cooling Inc. has a 6-month backlog of orders for its patented solar heating system. To meet this demand, management plans to expand production capacity by 25% with a $15 million investment in plant and machinery. The firm wants to maintain a 40% debt level in its capital structure. It also wants to maintain its past dividend policy of distributing 50% of last year's net income. In 2018, net income was $5 million. How much external equity must Coastal Carolina seek at the beginning of 2019 to expand capacity as desired? Assume that the firm uses only debt and common equity in its capital structure. Write out your answer completely. For example, 25 million should be entered as 25,000,000. Round your answer to the nearest cent.

$  

After a 2-for-1 stock split, Tyler Company paid a dividend of $1.9 per new share, which represents a 12% increase over last year's pre-split dividend. What was last year's dividend per share? Round your answer to the nearest cent.

$  
Answer #1

1). Retained earnings = Net income * (1 - Payout ratio)

= $5,000,000 * (1 - 0.50) = $2,500,000

Total equity required = New investment * (1 - Debt-to-capital ratio)

= $15,000,000(1 - 0.40) = $9,000,000

New external equity needed = Total Equity Required - Retained Earnings

= $9,000,000 - $2,500,000 = $6,500,000

2). DPS after split = $1.9.

Equivalent pre-split dividend = $1.9 * 2 = $3.8.

New equivalent dividend = Last year’s dividend * (1 + g)

$3.8 = Last year’s dividend * (1 + 0.12)

Last year’s dividend = $3.8/1.12 = $3.39